Let C be the curve defined by the parametric equations:

x = 4t2 + 4t

y = 2cos(t) + 3

t ∈ [0,π]


a) Find the value of on the indicated domain when the x-coordinate of the curve is 3. Write your answer as a fraction

t =


b) Find the value of on the indicated domain when the y-coordinate of the curve is 2. Round your answer to 3 decimal places 


t =


c) Find the y-intercept of the curve.

The y-intercept is at ( 0,_)




1
Expert's answer
2021-03-30T07:42:05-0400

а) The x-coordinate of the curve is 3, if:


"4t^2+4t=3""4t^2+4t-3=0""D=4^2-4\cdot4\cdot(-3)=16+48=64""t_1=\frac{-4-\sqrt{64}}{2\cdot4}=\frac{-4-8}{8}=-1.5""t_2=\frac{-4+\sqrt{64}}{2\cdot4}=\frac{-4+8}{8}=0.5"

Since "t_1 \notin [0; \pi]", then x =3 at t=0.5


b) The y-coordinate of the curve is 2, if:


"2cos(t) + 3=2""2cos(t) =-1""cos(t) = -\frac{1}{2}"

This equation has one solution on the interval "[0; \pi]": "t = \frac{2\pi}{3}"

c) The y-intercept is an (x,y) point with x=0. Hence,


"4t^2+4t=0""4t(t+1)=0""\left[ 
 \begin{gathered} 
 t=0\\t+1=0
 \end{gathered} 
\right.""\left[ 
 \begin{gathered} 
 t=0\\t=-1
 \end{gathered} 
\right."

Since "t=-1 \notin [0; \pi]", then x=0 at t=0. Find y at t=0:


"y=2cos(0)+3=2+3=5"

Hence, the  y-intercept is at (0,5)